2. (10 pts) You borrow $40,000 from the local bank at 5% APR, compounded monthly on a 6-year loan but you want to pay for the loan with payments every four months. (a) What is your effective interest on your 4-month payment? (b) What is your payment made every four months? (c) What is the APY for your loan with your payments??
Answer #1

Paste B TU WA Font Format Painter Clipboard G11 I B C D 40000 00 vouw Loan APR Years 5% 6 1.68% 2592.94 9 10 11 12 5.12%

LIPUUalu IUNIL HIITICIIL D7 A в с т р nm + in Loan APR Years 40000 0.05 6 7 8 =(1+C5/12)^4-1 =PMT(C8,C6*3,-C4) a so 12 13 =(1